# CVVNet: A Cross-Vertical-View Network for Gait Recognition

### v2

@@ Abstract
Gait recognition enables contact-free, long-range person identification that is robust to clothing variations and non-cooperative scenarios. While existing methods perform well in controlled environments, they struggle with cross-vertical view scenarios where elevation changes induce severe perspective distortions and self-occlusions. Our analysis reveals conventional CNN-based approaches relying on single-scale receptive fields fail to capture hierarchical frequency features, while standard self-attention mechanisms tend to over-smooth critical high-frequency edge patterns. To tackle this challenge, we propose Cross-Vertical-View Network(CVVNet), a frequency aggregation architecture specifically designed for robust cross-vertical-view gait recognition. CVVNet employs a Multi-Scale Attention Gated Aggregation (MSAGA) module which consists of a High-Low Frequency Extraction module (HLFE) and a Dynamic Gated Aggregation (DGA) mechanism. Specifically, HLFE adopts parallel multi-scale convolution/max-pooling path and self-attention path as high- and low-frequency mixers for effective multi-frequency feature extraction from input silhouettes, while DGA is introduced to adaptively adjust the fusion ratio of frequency and spatial features. The integration of HLFE and DGA enables CVVNet to effectively handle distortions from view changes, significantly improving the recognition robustness across different vertical views. Experimental results show that our CVVNet achieves state-of-the-art performance, with 8.6% improvement on DroneGait and 2% on Gait3D compared with the best existing methods.
@@ 4.1 Datasets and Implementation Details
The optimization setup uses AdamW with base learning rate $1\times 10^{-4}$ and weight decay $0.05$, coupled with OneCycleLR scheduler configured with maximum learning rate $6\times 10^{-4}$, total training steps $80,\!000$, and warm-up phase covering $6\%$ of the cycle. In particular, identical optimizer and scheduler configurations are strictly maintained for both Gait3D and DroneGait datasets to ensure methodological consistency.
@@ 4.2 Results and Analysis
Under matched-view conditions, CVVNet maintains state-of-the-art performance, achieving a near-perfect accuracy of 99.6% NM, 99.7% BG and 99.9% CL at low vertical views. For mid-view recognition, the method reaches 97.3% NM accuracy, surpassing GaitBase by 0.5%, while marginally trailing the SOTA methods by 0.2% and 0.1% in the BG and CL metrics, respectively. In high vertical view scenarios, our method leads SOTA methods by 1.2% in NM and 0.5% in CL, and trails by only 0.3% in BG. These experimental results collectively validate CVVNet’s recognition robustness and consistent superiority in both ideal and real-world surveillance scenarios.
@@ 4.3 Ablation studies
To quantify the contributions of the HLFE and DGA components in MSAGA, we conduct ablation studies on the DroneGait dataset under low-to-high vertical view conditions. As shown in Tab 5, HLFE consistently surpasses the P3D baseline across aggregation strategies, achieving its largest gain of 3.1% over P3D-DGA in cross-view closed-loop scenarios with 36.7% accuracy.
The DGA mechanism demonstrates critical advantages over static fusion methods. Replacing Add/Concat operations with dynamic gating improves accuracy by 1.3-2.9%, with HLFE-DGA reaching 34.3% accuracy—3.0% higher than HLFE-Concat. This confirms DGA’s ability to contextually reweight multi-frequency features rather than applying fixed fusion rules. Notably, the combined HLFE-DGA framework delivers a 6.7% closed-loop accuracy improvement over P3D-Add, where HLFE contributes 2.1% through enhanced low-frequency rhythm modeling while DGA provides 3.0% via adaptive fusion. These results jointly validate that effective cross-view recognition requires both multi-scale feature decomposition and intelligent pattern integration.

### v3

@@ Abstract
Gait recognition enables contact-free, long-range person identification that is robust to clothing variations and non-cooperative scenarios. While existing methods perform well in controlled environments, they struggle with cross-vertical view scenarios where elevation changes induce severe perspective distortions and self-occlusions. Our analysis reveals conventional CNN-based approaches relying on single-scale receptive fields fail to capture hierarchical frequency features, while standard self-attention mechanisms tend to over-smooth critical high-frequency edge patterns. To tackle this challenge, we propose Cross-Vertical-View Network(CVVNet), a frequency aggregation architecture specifically designed for robust cross-vertical-view gait recognition. CVVNet employs a Multi-Scale Attention Gated Aggregation (MSAGA) module which consists of a High-Low Frequency Extraction module (HLFE) and a Dynamic Gated Aggregation (DGA) mechanism. Specifically, HLFE adopts parallel multi-scale convolution/max-pooling path and self-attention path as high- and low-frequency mixers for effective multi-frequency feature extraction from input silhouettes, while DGA is introduced to adaptively adjust the fusion ratio of frequency and spatial features. The integration of HLFE and DGA enables CVVNet to effectively handle distortions from view changes, significantly improving the recognition robustness across different vertical views. Experimental results show that our CVVNet achieves state-of-the-art performance on both datasets, delivering an 8.6% improvement on DroneGait while retaining the leading position on Gait3D.
@@ 4.1 Datasets and Implementation Details
The optimization setup uses AdamW with base learning rate $1\times 10^{-4}$ and weight decay $0.05$, coupled with OneCycleLR scheduler configured with maximum learning rate $6\times 10^{-4}$, total training steps 80,00080,\!000, and warm-up phase covering $6\%$ of the cycle. In particular, identical optimizer and scheduler configurations are strictly maintained for both Gait3D and DroneGait datasets to ensure methodological consistency.
@@ 4.2 Results and Analysis
Under matched-view conditions, CVVNet maintains state-of-the-art performance, achieving a near-perfect accuracy of 99.6% NM, 99.7% BG and 99.9% CL at low vertical views. For mid-view recognition, the method reaches 97.3% NM accuracy, surpassing GaitBase by 0.5%, while marginally trailing the SOTA methods by 0.1% and 0.1% in the BG and CL metrics, respectively. In high vertical view scenarios, our method leads SOTA methods by 1.2% in NM and 0.5% in CL, and trails by only 0.3% in BG. These experimental results collectively validate CVVNet’s recognition robustness and consistent superiority in both ideal and real-world surveillance scenarios.
@@ 4.3 Ablation studies
To quantify the contributions of the HLFE and DGA components in MSAGA, we conduct ablation studies on the DroneGait dataset under low-to-high vertical view conditions. As shown in Tab 5, HLFE consistently demonstrated superior performance compared to the P3D baseline. Across all aggregation strategies, configurations employing HLFE consistently outperformed their respective P3D counterparts. Notably, in the most challenging CL cross-view scenario, the HLFE-DGA model achieved a Rank-1 accuracy of 36.7%, representing a 3.1% improvement over P3D-DGA (33.6%). This highlights the effectiveness of HLFE in extracting and encoding crucial low-frequency gait information, particularly under complex viewpoint variations.
The DGA mechanism demonstrates critical advantages over static fusion methods. Replacing traditional fusion methods with DGA consistently resulted in accuracy improvements ranging from 1.3% to 3.1% across different feature extractors (P3D or HLFE) and various view scenarios. For instance, within the HLFE series, DGA improved accuracy by 1.3% over Concat (from 33.0% to 34.3%) and by 2.9% over Add (from 31.4% to 34.3%) in the NM scenario. This result strongly confirms DGA’s ability to dynamically weight multi-frequency features based on context, rather than applying fixed fusion rules, thereby enabling more effective information integration.